OpenStudy (anonymous):

MEDAL WILL BE AWARDED Which graph best represents the function f(x) = 3(1.5)x? graph of increasing exponential function going through point 0, 2 graph of increasing exponential function going through point 0, 3 graph of increasing exponential function going through point 0, 1 graph of increasing exponential function going through point 0, 4

OpenStudy (anonymous):

@arabpride @BlossomCake @gold_grace123 @Kainui CAN ANYONE HELP?!?!?

OpenStudy (anonymous):

Medal will be awarded

OpenStudy (anonymous):

I would love to help, but I only just started learning about Functions. Very sorry, I hope someone can help you with your question!

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@jagr2713 @blackops2luvr @KJ4UTS @LeeEtchison Anyone help?

OpenStudy (anonymous):

@mathstudent55 @Mehek14 @Abmon98

Mehek (mehek14):

is it f(x) = 3(1.5)x or f(x) = 3(1.5)^x?

OpenStudy (anonymous):

im not sure, i dont know anything about this question

Mehek (mehek14):

well where did u get this question from? make sure and see which one it is

OpenStudy (anonymous):

its from flvs

Mehek (mehek14):

so check which one

OpenStudy (anonymous):

its from algebra 1 segment 1 exam part 2 question 1

OpenStudy (anonymous):

@TwilightRain1 @ichliebedich can you guys try to help?

Mehek (mehek14):

oh im doing that too

OpenStudy (anonymous):

really?!

Mehek (mehek14):

yea who's class r u in?

OpenStudy (anonymous):

mrs yelvington

Mehek (mehek14):

oh im not in that one

OpenStudy (anonymous):

oh...

OpenStudy (anonymous):

wait, so do you know the answer?

OpenStudy (anonymous):

@KrazyKricket

OpenStudy (anonymous):

can you help me, i will give medal?

Nnesha (nnesha):

use that points one by one and substitute that into y and x if you get both side equal then that is your answer

OpenStudy (anonymous):

so, would it be c?

Nnesha (nnesha):

and yes it is \[\huge f(x)= 3(1.5)^x\]

OpenStudy (anonymous):

so its c?

Nnesha (nnesha):

nope :(

OpenStudy (anonymous):

oh

Nnesha (nnesha):

replace x by 0 and remember when you have something to the power of 0 its just turn to 1

OpenStudy (anonymous):

ok wait, so its B?

Nnesha (nnesha):

are you sure ??

OpenStudy (anonymous):

yes

Nnesha (nnesha):

yes :)

OpenStudy (anonymous):

thank you so much, i will give you a medal :)

Nnesha (nnesha):

my pleasure :)